OpenStudy (anonymous):

What is the missing expression x8 + 3x5 = x5(_____)

OpenStudy (anonymous):

Please use the ^ symbol to indicate exponents... Please...

OpenStudy (anonymous):

x^8 + 3x^5 = x^5(_____) This?

OpenStudy (anonymous):

i forgot i remember u i was about to fix it sry

OpenStudy (anonymous):

\[x^{8-5} + 3x^{5-5} = x^3 + 3x^0 = x^3+3\] See how this works now? :-)

OpenStudy (anonymous):

You are factoring out a x^5

OpenStudy (anonymous):

dividing exponents, you subtract the power #'s

OpenStudy (anonymous):

and anything to the 0 power is ALWAYS 1 (because mathematicians said so), even \( 0^0 \)

OpenStudy (anonymous):

x^3+3x=x^5???

OpenStudy (anonymous):

no... \[x^8 + 3x^5=x^5(x^3+3) \]

OpenStudy (anonymous):

look at the numbers on the exponents ;-)

OpenStudy (anonymous):

And re-read what I wrote above, it should click. If you're factoring stuff up... \[ax+bx \rightarrow \frac{ax+bx}{x}=x(a+b)\]